OpenStudy (anonymous):

When the quadratic equation y = (x - 5)(2x + 3) is written in standard form, what is the value of the coefficient "b"?

OpenStudy (anonymous):

are you familiar with the FOIL method?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

ok, when you FOIL it out, the standard form is: ax^2+bx+c

OpenStudy (anonymous):

so it will be the coefficient of the x variable

OpenStudy (anonymous):

the coefficient of x with degree 1

OpenStudy (anonymous):

ay+bx+c=0 --> Standard Form y=(x-5)(2x+3) =2x^2+3x-10x-15 =2x^2-7x-15

OpenStudy (anonymous):

thanks you guys